OpenStudy (anonymous):

PLEASE HELP!!!! A point H on a segment with endpoints B(3, −1) and Z(12, 5) partitions the segment in a 5:1 ratio. Find H. You must show all work to receive credit.

OpenStudy (misty1212):

HI!!

OpenStudy (misty1212):

this looks hard, but i bet we can do it

OpenStudy (anonymous):

i totally dont get it lol

OpenStudy (misty1212):

yeah lets think for second and see what we can come up with

imqwerty (imqwerty):

u have a formula to find this :)

OpenStudy (misty1212):

first off that means we are going to break it in to 6 equal parts, and take 5 of them

OpenStudy (misty1212):



imqwerty (imqwerty):

the cordinates of point P(x,y) which lies on segment AB with end points A(a,b) and B(c,d) and divides it in the ration m:n so u have - x=(mc+na)/(m+n) and y=(md+nb)/(m+n)

OpenStudy (misty1212):

yeah i don't get that formula from 3 to 12 is 9 units one sixth of that is \(\frac{9}{3}=\frac{3}{2}\)

imqwerty (imqwerty):